OpenStudy (anonymous):

The vertex of this parabola is at (-3, -1). When the y-value is 0, the x-value is 4. What is the coefficient of the squared term in the parabola's equation?

OpenStudy (anonymous):

help me please

OpenStudy (zale101):

Apparently, you have 2 given points the vertex and the x-intercept the vertex is (-3,-1) the x-intercept is (4,0) clear?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i dont understand any

OpenStudy (zale101):

what part of what i just said didn't make sense to you?

OpenStudy (anonymous):

well i dont understand how to get the answer

OpenStudy (zale101):

i know you dont understand on how to find the answer, but what about the description i made, did it make sense ?

OpenStudy (anonymous):

yes

OpenStudy (zale101):

The equation of a parabola in a vertex form \[y = a(x - h)^2 + k\] vertex (h,k)=(-3,-1) x intercept x,y=(4,0) tat means the only variable left is a, solve for it

OpenStudy (zale101):

Sorry about the confusion earlier, i should have used y=a(x-h)^2+k since you already mentioned it to me.

OpenStudy (dan815):

Here is a summary of the 3 main points

OpenStudy (dan815):

So you are given 2 points, and one of them is the vertex, at the vertex f'(x) = 0 f ' (-3)=0 (1) f(-3) =1 (2) f(4)=0 (3) So you have these 3 pieces of information

OpenStudy (dan815):

A parabola is a quadratic of the form y=ax^2+bx+c you have 3 unknowns and you can form 3 linearly independent equations y'(-3)=0 y(-3)=1 y(4)=0 solve for a,b,c

OpenStudy (dan815):

that 1 should be -1, fix that and carry on